OpenStudy (anonymous):

A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue: Write two functions to model this situation, and graph those two functions on the same coordinate grid. During what year in the future will the park have approximately the same number of pine and oak trees?

OpenStudy (anonymous):

@musicmaster @rockstar0765 @darkb0nebeauty @TheGaMer007 @nuccioreggie can yall answer A and b FOR ME PLEASE , ILL GIVE BEST ANSWER, FRIEND. AND TESTIMONIAL @sleepyjess @sammixboo @dan815

OpenStudy (anonymous):

@radar @tywower

OpenStudy (anonymous):

@peachpi

OpenStudy (welshfella):

the formula for number of pines after n years is N = 800( 1- 0.050)^n

OpenStudy (welshfella):

can you work out what it would be for the oaks?

OpenStudy (anonymous):

ummm idk, ill tyr

OpenStudy (anonymous):

try

OpenStudy (welshfella):

14% = 0.14

OpenStudy (anonymous):

wats that for? ( the .14)?

OpenStudy (welshfella):

look at the function for the pines the 0.05 come from the 5% this time its 0.14 but the number of trees are increasing so its + 0.14

OpenStudy (welshfella):

* OH sorry its 0.15 ( 15%)

OpenStudy (anonymous):

N = 800( 1- 0.15)^n right?

OpenStudy (welshfella):

no - the number is INCREASING not decreasing

OpenStudy (anonymous):

N = 800( 1+0.15)^n

OpenStudy (welshfella):

lol I'll do that again N = 800(0.95)^n and N = 800(1.14)^n

OpenStudy (welshfella):

- these are the same as the others but more concise

OpenStudy (anonymous):

are those the 2 functions to the first question?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ohh lol ok, thankyou. wb the second question

OpenStudy (welshfella):

how do you think you answer the second part - when number of pines = number of oaks?

OpenStudy (welshfella):

when that happens the value N will be the same for oaks and pines

OpenStudy (anonymous):

is it the 14th year?

OpenStudy (anonymous):

o nevermind sorry

OpenStudy (welshfella):

how did you get that?

OpenStudy (welshfella):

thats 0.15 not 0.14 my last post - i seem to have a fixation with 14!!!

OpenStudy (anonymous):

lol ok, thankyou!

OpenStudy (welshfella):

i've made another error for the oak trees the number is 50 not 800 so that the second function is 50(1.15)^n sorry